OpenStudy (anonymous):

convert between standard & vertex form y=-2x^2-24x-79

OpenStudy (anonymous):

Are you familiar with the process of completing the square?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So lets do that here: \[y = -2(x^2 + 12x +0 ) - 79\]\[y = -2(x^2 + 12x + 36 - 36) - 79\]\[y=-2(x^2 + 12x + 36) - 2(-36) - 79\]\[y=-2(x^2 + 12x +36) + 72 - 79\]\[y = -2(x^2 + 12x + 36) -7\]

OpenStudy (anonymous):

and (x^2 + 12x + 36) = (x + 6)^2

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

6 and 36

OpenStudy (anonymous):

6 and 36 is not the vertex form of the function.

OpenStudy (anonymous):

Isn't that what you're trying to find? the vertex form?

OpenStudy (anonymous):

yes